Find all real numbers A,B such that the function f (x) = A(2^x) + B, satisfies the conditions f`(1) = 2 and
integration 0 to 3 f(x)dx = 7.

Grade:11

1 Answers

Sher Mohammad IIT Delhi
9 years ago
f'(x)=2a=2
a=1
f(x)=2x+b
integral f(x)=x^2+bx
putting limits
(9+3b)=7
3b=-2
b=-2/3

sher mohammad
b.tech, iit delhi